how do i solve 4x+3y=14 .... using the addition method
9x-2y=14

Answer by best82(13)

Is this two separate equations??
4x+3y=14
9x-2y=14
In addition method, you want to eliminate one variable. Let's get rid of the y.
You must get the y in both equations to be the same thing. The LCM is 6. Then multiply each equation by a certain number, such that the number in front of y is 6. One must be negative, since -+ (+)= cancelled.
-2(4x+3y)=2(14)
3(9x-2y)=3(14)
-8x-6y=28
27x-6y=42
Get rid of the y.
19x=72
x=4
y=11